### v2

# Parameter estimation for partially observed second-order diffusion processes

Abstract.
Estimating parameters of a diffusion process given continuous-time observations of the process via maximum likelihood approaches or, online, via stochastic gradient descent or Kalman filter formulations constitutes a well-established research area. It has also been established previously that these techniques are, in general, not robust to perturbations in the data in the form of temporal correlations of the driving noise. While the subject is relatively well understood and appropriate modifications have been suggested in the context of multi-scale diffusion processes and their reduced model equations, we consider here an alternative but related setting where a diffusion process in positions and velocities is only observed via its positions. In this note, we propose a simple modification to standard stochastic gradient descent and Kalman filter formulations, which eliminates the arising systematic estimation biases. The modification can be extended to standard maximum likelihood approaches and avoids computation of previously proposed correction terms.
1. Introduction
We consider the problem of online learning parameters $\theta\in\mathbb{R}^{d_{\theta}}$ of second-order stochastic differential equations (SDEs)
(1a) ${\rm d}U_{t}$ $=f(X_{t},U_{t}){\rm d}t+F(X_{t},U_{t})\theta{\rm d}t+\sqrt{\sigma}{\rm d}W_{t},$
(1b) ${\rm d}X_{t}$ $=U_{t}{\rm d}t$
from incoming observations $X_{t}^{\dagger}\in\mathbb{R}^{d}$, $t\geq 0$, of an underlying reference process. Here $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $F:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d_{\theta}}$ are given functions, $\sigma>0$ is the diffusion constant, and $W_{t}$ denotes $d$-dimensional Brownian motion. The data $X_{t}^{\dagger}$ is being generated from a second-order SDE of the same type, i.e.,
(2a) ${\rm d}U_{t}^{\dagger}$ $=g(X_{t}^{\dagger},U_{t}^{\dagger}){\rm d}t+\sqrt{\sigma}{\rm d}W_{t}^{\dagger},$
(2b) ${\rm d}X_{t}^{\dagger}$ $=U_{t}^{\dagger}{\rm d}t$
with the drift $g(x,u)$ being unknown.
Note that for $f,F,g$ independent of position, the system corresponds to a first-order diffusion process where only the integral of the process is being observed.
We assume that $g(x,u)$ is such that the reference process $(U_{t}^{\dagger},X_{t}^{\dagger})$ is ergodic with invariant distribution $\pi$. The optimal parameter $\theta_{\ast}$ can then be determined by minimizing the ergodic least squares functional
(3) $\mathcal{L}(\theta)=\frac{1}{2}\mathbb{E}_{\pi}\left[\|g(X,U)-f(X,U)-F(X,U)\theta\|^{2}\right],$
which has a unique minimizer provided the Fisher information matrix $\mathbb{E}_{\pi}[F(X,U)^{T}F(X,U)]$ has full rank. Moreover, we assume that the unique minimizer satisfies
(4) $g(x,u)=f(x,u)+F(x,u)\theta_{\ast}\,.$
However, the gradient of $\mathcal{L}(\theta)$ cannot be computed since $g(x,u)$ is unknown. Therefore we investigate online estimation techniques in this paper based on observed positions $X_{t}^{\dagger}$.
While standard online estimation techniques, such as stochastic gradient descent (SGD) [21] and Kalman filtering [14, 19], can be applied in case both ($X_{t}^{\dagger},U_{t}^{\dagger})$ from (2) are being observed, the estimation problem becomes more challenging in case only positions $X_{t}^{\dagger}$ are being observed and velocities need to be estimated using finite differencing [18, 6], such as,
(5) $\tilde{U}_{n+1/2}:=\frac{X_{t_{n+1}}^{\dagger}-X_{t_{n}}^{\dagger}}{\tau}$
with $\tau>0$ denoting the sampling interval, $t_{n}=\tau n$, for all $n\geq 0$.
As demonstrated previously [10, 12, 11], naive application of standard maximum likelihood-based (MLE) estimators [13, 16] leads to a systematic bias. In order to correct for the arising estimation bias, appropriate corrections terms to the MLE estimator have been proposed in [12, 6, 3, 11]. An alternative approach has been to treat the continuous time velocity process as a latent variable and calculate the MLE estimator using an expectation-maximization algorithm [4]. Viewing the SDEs as a partially observed system, the likelihood of $\theta$ could also be approximated using an extended Kalman filter [20]. Related phenomena have been investigated in the context of multi-scale diffusion [2, 15, 5], which again can be corrected via an appropriately modified MLE estimator [9].
While these correction terms can be applied to online learning, we follow an alternative approach in this note, which has been motivated by the related work [1, 17]. Our approach can easily be implemented both for SGD [21] as well as Kalman filter [14, 19] formulations.
As demonstrated in the next section by means of a simple example, a discretized SGD of the form
(6a) $\theta_{n+1}$ $:=\theta_{n}+\alpha_{t}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})^{\rm T}\Delta I_{n}(\theta_{n}),$
(6b) $\Delta I_{n}(\theta)$ $:=(\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2})-\left(f(X_{t_{n}}^{\dagger},\tilde{U}_{n})+F(X_{t_{n}}^{\dagger},\tilde{U}_{n})\theta\right)\tau,$
with $U_{t_{n+1/2}}$ being replaced by its estimated $\tilde{U}_{n+1/2}$ etc. and the symmetric approximation
(7) $\tilde{U}_{n}:=\frac{1}{2}\left(\tilde{U}_{n+1/2}+\tilde{U}_{n-1/2}\right)=\frac{X_{t_{n+1}}^{\dagger}-X_{t_{n-1}}^{\dagger}}{2\tau}$
[…]
(8) $\alpha_{t}=c_{1}/(c_{2}+t)$
denotes the learning rate for appropriate constants $c_{i}>0$, $i=1,2$. While the centered approximation in (LABEL:eq:U_symmetric_is_the_best_approximation_to_\$U_n\$_from_the_position_measurements_at_the_nearest_neighbor_time_points_\$X_t_n-1,_X_t_n,X_t_n+1\$_alternative_approximations_to_the_velocity_at_\$t_n\$_could_also_be_employed.)
The discussion from the following section will lead to an appropriate modification of (6), which removes the undesirable bias in the SGD estimator (6). See Section 3.
We extend the proposed methodology to Kalman filter based estimation in Section 4. The unbiasedness of the corrected SGD approach is demonstrated for a nonlinear second-order diffusion process in Section 5. The note closes with comments on further directions for research.
2. Motivating example
In order to demonstrate the difficulties the standard SGD implementation (6) encounters, we consider velocity data $U_{t}^{\dagger}$ being generated by the linear OU process
(9) ${\rm d}U_{t}^{\dagger}=-\theta_{\ast}U_{t}^{\dagger}{\rm d}t+\sqrt{\sigma}{\rm d}W_{t}^{\dagger}$
for given diffusion constant $\sigma>0$ and unknown drift parameter $\theta_{\ast}\geq 0$.
We will set $\theta_{\ast}=0$ for some of the analysis to be conducted later in this section, which implies that solutions of (9) satisfy
(10) $U_{t}^{\dagger}=U_{s}^{\dagger}+\sqrt{\sigma}W_{t-s}^{\dagger}$
for any $t\geq s$ and $s\geq 0$ and, consequently,
(11) $X_{t}^{\dagger}=X_{s}^{\dagger}+(t-s)U_{s}^{\dagger}+\sqrt{\sigma}\int_{s}^{t}W^{\dagger}_{r-s}{\rm d}r.$
Here, we have introduced the shorthand $W_{t-s}^{\dagger}=W_{t}^{\dagger}-W_{s}^{\dagger}$.
We first recall some of the properties of SGD when a solution of (9) has in fact been fully observed. An application of SGD [21] to this problem yields
(12a) ${\rm d}\theta_{t}$ $=-\alpha_{t}U_{t}^{\dagger}{\rm d}I_{t}=-\alpha_{t}U_{t}^{\dagger}\left({\rm d}U_{t}^{\dagger}+\theta_{t}U_{t}^{\dagger}{\rm d}t\right)$
(12b) $=-\alpha_{t}U_{t}^{\dagger}\left((\theta_{t}-\theta_{\ast})U_{t}^{\dagger}{\rm d}t+\sqrt{\sigma}{\rm d}W_{t}^{\dagger}\right).$
We call
(13) $I_{t}=\int_{0}^{t}\left({\rm d}U_{s}^{\dagger}+\theta_{s}U_{s}^{\dagger}{\rm d}s\right)=\int_{0}^{t}\left((\theta_{s}-\theta_{\ast})U_{s}^{\dagger}{\rm d}s+\sqrt{\sigma}{\rm d}W_{s}^{\dagger}\right)$
the innovation and $\alpha_{t}U_{t}^{\dagger}$ the gain. It is obvious that $I_{t}$ becomes a martingale at the optimal parameter choice $\theta_{s}=\theta_{\ast}$, $s\in[0,t]$.
It has been shown that $\theta_{t}\to\theta_{\ast}$ as $t\to\infty$ [21]. In our context, a key observation is that the integral
(14) $\int_{0}^{t}\alpha_{s}U_{s}^{\dagger}\left({\rm d}U_{s}^{\dagger}+\theta_{t}U_{s}^{\dagger}{\rm d}s\right)$
also becomes a martingale for any $t>0$ if and only if $\theta_{s}=\theta_{\ast}$ for all $s\in[0,T]$.
In terms of numerical implementations of the SGD formulation (12), we consider the discrete time SGD method
(15) $\theta_{n+1}:=\theta_{n}-\alpha_{t_{n}}U_{t_{n}}^{\dagger}\left(U_{t_{n+1}}^{\dagger}-U_{t_{n}}^{\dagger}+\frac{\theta_{n}}{2}\left(U_{t_{n+1}}^{\dagger}+U_{t_{n}}^{\dagger}\right)\tau\right).$
This discrete-time SGD method still implies that $\theta_{n}$ converges to $\theta_{\ast}$ as $n\to\infty$ and $\tau\to 0$. More precisely, any finite sum
(16) $\sum_{n=0}^{N-1}U_{t_{n}}^{\dagger}\left(U_{t_{n+1}}^{\dagger}-U_{t_{n}}^{\dagger}+\frac{\theta_{n}}{2}\left(U_{t_{n+1}}^{\dagger}+U_{t_{n}}^{\dagger}\right)\tau\right)$
becomes a martingale as $\tau\to 0$ for $T=N\tau$ fixed if and only if $\theta_{n}=\theta_{\ast}$.
However, the focus of this note is on the case when only positions $X_{t}^{\dagger}$, $t\geq 0$, can be observed and we recall the velocity approximations (5), where the sampling interval $\tau>0$ can be made arbitrarily small. We note that (5) can also be viewed as a local time average, i.e.,
(17) $\tilde{U}_{n+1/2}=\frac{1}{\tau}\int_{t_{n}}^{t_{n+1}}U_{t}^{\dagger}{\rm d}t,$
which provides a link to the related work [12, 8].
Using (7) instead of the true $U_{t_{n}}^{\dagger}$ etc.  and symmetrizing the gain leads to the previously stated (6), which reduces here to
(18a) $\theta_{n+1}$ $:=\theta_{n}-\alpha_{t_{n}}\tilde{U}_{n}\left(\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2}+\theta_{n}\tilde{U}_{n}\tau\right)$
(18b) $=\theta_{n}-\frac{\alpha_{t_{n}}}{2}\left(X_{t_{n+1}}^{\dagger}-X_{t_{n-1}}^{\dagger}\right)\left(\frac{X_{t_{n+1}}^{\dagger}-2X_{t_{n}}^{\dagger}+X_{t_{n-1}}^{\dagger}}{\tau^{2}}+\frac{\theta_{n}}{2}\frac{X_{t_{n+1}}^{\dagger}-X_{t_{n-1}}^{\dagger}}{\tau}\right).$
We note that, while the discrete innovations
(19) $I_{N}=\sum_{n=1}^{N-1}\left(\frac{X_{t_{n+1}}^{\dagger}-2X_{t_{n}}^{\dagger}+X_{t_{n-1}}^{\dagger}}{\tau^{2}}+\frac{\theta_{n}}{2}\frac{X_{t_{n+1}}^{\dagger}-X_{t_{n-1}}^{\dagger}}{\tau}\right)\tau$
still converge to a martingale as $\tau\to 0$ for $T=N\tau$ fixed provided $\theta_{n}=\theta_{\ast}$ for all $n\geq 0$, this is no longer true for
(20) $\frac{1}{2}\sum_{n=1}^{N-1}\left(X_{t_{n+1}}^{\dagger}-X_{t_{n-1}}^{\dagger}\right)\left(\frac{X_{t_{n+1}}^{\dagger}-2X_{t_{n}}^{\dagger}+X_{t_{n-1}}^{\dagger}}{\tau^{2}}+\frac{\theta_{n}}{2}\frac{X_{t_{n+1}}^{\dagger}-X_{t_{n-1}}^{\dagger}}{\tau}\right).$
This violation of the martingale property is perhaps not surprising since (20) does not resemble a discretized Itô-Integral like (16) but rather a discretized Stratonovitch integral [16]. We note that such an interpretation as a discretized stochastic integral concerns only the second-order differencing term in (20) and we may therefore set $\theta=\theta_{\ast}=0$ in the subsequent derivations. Recall that
(21) $\frac{X_{t_{n+1}}-X_{t_{n}}}{\tau}=\tilde{U}_{n+1/2}=U_{t_{n}}^{\dagger}+\frac{\sqrt{\sigma}}{\tau}\int_{t_{n}}^{t_{n+1}}W_{t-t_{n}}^{\dagger}{\rm d}t$
and we therefore obtain
(22a) $\frac{1}{2}\mathbb{E}\left[\left(\tilde{U}_{n+1/2}+\tilde{U}_{n-1/2}\right)\left(\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2}\right)\right]$ $=\frac{1}{2}\mathbb{E}\left[\tilde{U}_{n+1/2}^{2}-\tilde{U}_{n-1/2}^{2}\right]$
(22b) $=\frac{1}{2}\mathbb{E}\left[(U_{t_{n}}^{\dagger})^{2}-(U_{t_{n-1}}^{\dagger})^{2}\right]=\frac{\tau\sigma}{2},$
which implies an estimation bias induced by the discrete sum (20) even in the limit $\tau\to 0$.
In order to remove the resulting bias in the SGD formulation (18), we propose to simply shift the innovation term forward in time such that it does no longer overlap with the gain; thus regaining an Itô-type stochastic integral approximation. A related idea has been proposed and investigated in [17] in the context of multi-scale diffusion. This modification leads to the following implementation:
(23) $\theta_{n+1}:=\theta_{n}+\alpha_{t_{n}}\tilde{U}_{n}\left(\tilde{U}_{n+5/2}-\tilde{U}_{n+3/2}-\theta_{n}\tilde{U}_{n+2}\tau\right).$
Note that the innovation in the modified implementation (23) only contains positional data from time points that are larger than or equal to those in the gain.
Let us introduce the random variables
(24) $\tilde{\Xi}_{n}:=\frac{1}{\sqrt{\sigma\tau}}\left(\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2}\right)$
for $n\geq 1$, which is a rescaling of the innovation $\Delta I_{n}(\theta_{\ast})$, as defined in (6b), with $\theta_{\ast}$ again set to $\theta_{\ast}=0$. It is crucial to observe that $\tilde{\Xi}_{n+2}$ is correlated with both $\tilde{U}_{n+5/2}$ and $\tilde{U}_{n+3/2}$, while being independent of $\tilde{U}_{n}$. Hence, provided $\tilde{\Xi}_{n}$ is a martingale for all $n\geq 1$, the SGD scheme (23) will become unbiased in the limit $\tau\to 0$.
Hence let us analyze the statistical properties of $\tilde{\Xi}_{n}$. We find that
(25a) $\tilde{\Xi}_{n}$ $=\frac{1}{\sqrt{\sigma\tau}}\left(\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2}\right)$
(25b) $=\frac{1}{\sqrt{\sigma\tau}}\left(U_{t_{n}}^{\dagger}-U_{t_{n-1}}^{\dagger}\right)+\frac{1}{\sqrt{\tau}}\left(\frac{\int_{t_{n}}^{t_{n+1}}W_{t-t_{n}}^{\dagger}{\rm d}t-\int_{t_{n-1}}^{t_{n}}W_{t-t_{n-1}}^{\dagger}{\rm d}t}{\tau}\right)$
(25c) $=\frac{W_{t_{n}-t_{n-1}}^{\dagger}}{\sqrt{\tau}}+\frac{1}{\sqrt{\tau}}\left(\frac{\int_{t_{n}}^{t_{n+1}}W_{t-t_{n}}^{\dagger}{\rm d}t-\int_{t_{n-1}}^{t_{n}}W_{t-t_{n-1}}^{\dagger}{\rm d}t}{\tau}\right).$
It follows that $\tilde{\Xi}_{n}$ is Gaussian with mean zero and that $\tilde{\Xi}_{n+2}$ is indeed independent of $\tilde{U}_{n}$.
We recall the following two properties of integrated Brownian motion:
(26) $\mbox{var}\,\left(\int_{t_{n}}^{t_{n+1}}W_{t-t_{n}}^{\dagger}{\rm d}t\right)=\frac{1}{3}\tau^{3}$
and
(27) $\mathbb{E}\left[W_{t_{n}-t_{n-1}}^{\dagger}\left(\int_{t_{n-1}}^{t_{n}}W_{t-t_{n-1}}^{\dagger}{\rm d}t\right)\right]=\int_{t_{n-1}}^{t_{n}}\mathbb{E}\left[(W_{t-t_{n-1}}^{\dagger})^{2}\right]{\rm d}t=\frac{1}{2}\tau^{2}.$
Upon applying these results to (25c) when computing the variance of $\tilde{\Xi}_{n}$, it is found that
(28) $\mbox{var}\,(\tilde{\Xi}_{n})=\frac{2}{3}.$
In other words, we obtain that $\tilde{\Xi}_{n}\sim{\rm N}(0,2/3)$ in agreement with results from [12, 7]. This result carries over to $\theta_{\ast}\not=0$ since the arising additional terms vanish as $\tau\to 0$.
We finally formally investigate the time-continuous limit of the modified discrete-time SGD formulation (23). We first note that
(29) $\mathbb{E}[\tilde{\Xi}_{n}\tilde{\Xi}_{n+k}]=\left\{\begin{array}[]{cc}\frac{1}{6}&|k|=1\\ 0&|k|>1\end{array}\right.,$
which implies that
(30) $Z_{t}\mathrel{\mathop{:}}=\begin{cases}0&\text{if}\quad t<\tau\\ \sum_{k=1}^{\lfloor t/\tau\rfloor}\sqrt{\tau}\,\tilde{\Xi}_{k}&\text{else}\end{cases}$
converges to standard Brownian motion in the limit $\tau\to 0$. When using this result in the SGD formulation (23) for general $\theta_{\ast}$, we obtain
(31a) $\theta_{n+1}$ $\approx\theta_{n}-\alpha_{t_{n}}\tilde{U}_{n}\left((\theta_{n}-\theta_{\ast})\tilde{U}_{n+2}\tau+\sqrt{\sigma\tau}\,\tilde{\Xi}_{n+2}\right)$
(31b) $\approx\theta_{n}-\alpha_{t_{n}}\tilde{U}_{n}\left((\theta_{n}-\theta_{\ast})\tilde{U}_{n}\tau+\sqrt{\sigma\tau}\,\tilde{\Xi}_{n+2}\right)$
and the limit $\tau\to 0$ leads formally to the time-continuous formulation
(32) $\dot{\theta}_{t}=-\alpha_{t}U_{t}^{\dagger}\left((\theta-\theta_{\ast})U_{t}^{\dagger}+\sqrt{\sigma}\dot{W}_{t}^{\dagger}\right),$
which coincides with (12b).
Remark: Independent of the drift estimation, it is well known that the diffusion term can be estimated from the variability of a process. In line with previous results [7, 11, 12], our above calculations and equation (28) imply the following estimator for the diffusion constant $\sigma$ from the quadratic variation of the approximated velocities $\tilde{U}_{n+1/2}$:
(33) $\sigma_{N}:=\frac{3}{2\tau N}\sum_{n=1}^{N}\|\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2}\|^{2}.$
3. Unbiased SGD
The analysis from Section 2 leads to the following modification of the discrete-time SGD implementation (6):
(34) $\theta_{n+1}=\theta_{n}+\alpha_{t_{n}}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})^{\rm T}\Delta I_{n+2}(\theta_{n}),$
for $n\geq 1$ with the innovations $\Delta I_{n}(\theta)$ defined by (6b) and the learning rate $\alpha_{t_{n}}$ by (8) with $t=t_{n}=n\tau$. We recall that the key point of shifting the innovation forward in time is to retain an Itô-type approximation in the gain times innovation stochastic integral approximation. This is the property that maintains the desired unbiasedness of the resulting sequential learning scheme in the limit $\tau\to 0$.
We note that unbiased formulations are not unique since only the shifting of the second-order differecing term in the innovation is relevant for consistency. For example, (34) could be replaced by
(35) $\theta_{n+1}=\theta_{n}+\alpha_{t_{n}}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})^{\rm T}\left(\tilde{U}_{n+5/2}-\tilde{U}_{n+3/2}-\left(f(X_{t_{n}}^{\dagger},\tilde{U}_{n})+F(X_{t_{n}}^{\dagger},\tilde{U}_{n})\theta_{n}\right)\tau\right).$
This freedom is also utilized when formulating the modified MLE estimator (39) further below in Section 6.
4. Unbiased Kalman filter
The SGD method (34) provides an online point estimator for the parameter $\theta$. However, it is also possible to implement an online Bayesian estimator using the Kalman filter [14, 19]. The key idea is to treat the unknown parameter as a Gaussian random variable $\Theta_{n}$ with mean $m_{n}$ and covariance matrix $\Sigma_{n}$ and with trivial dynamics $\Theta_{n+1}=\Theta_{n}$, $n\geq 0$ [20]. The evolution equation for the mean and the covariance matrix follow the standard Kalman update formulas:
(36a) $m_{n+1}$ $:=m_{n}+K_{n}\frac{\Delta I_{n+2}(m_{n})}{\tau},$
(36b) $\Sigma_{n+1}$ $:=\Sigma_{n}-K_{n}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})\Sigma_{n}$
with Kalman gain
(37) $K_{n}:=\Sigma_{n}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})^{\rm T}\left(F(X_{t_{n}}^{\dagger},\tilde{U}_{n})^{\rm T}\Sigma_{n}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})+\frac{\sigma}{\tau}I\right)^{-1}$
for $n\geq 1$. The iteration is initialized from the Gaussian prior $\Theta_{1}\sim{\rm N}(m_{\rm prior},\Sigma_{\rm prior})$.
In contrast to the SGD formulation (34), the learning rate $\alpha_{t_{n}}$ no longer appears and has instead been submerged implicitly into the Kalman gain $K_{n}$. Indeed, the singular values of the covariance matrix $\Sigma_{n}$ decay with rate $1/t_{n}$ under the already stated ergodicity assumption on (2) and the non-singularity of the Fisher information matrix $\mathbb{E}_{\pi}\left[F(X,U)^{\rm T}F(X,U)\right]$. If the required diffusion constant $\sigma$ is unknown, it can either be estimated via (33) or, alternatively, be treated as part of the prior covariance matrix $\Sigma_{\rm prior}$ and then set to $\sigma=1$ in the Kalman gain (37).
5. Numerical example
We consider the nonlinear second-order SDE
(38a) ${\rm d}U_{t}$ $=-X_{t}{\rm d}t+U_{t}{\rm d}t-\theta U_{t}^{3}{\rm d}t+\sqrt{2}{\rm d}W_{t},$
(38b) ${\rm d}X_{t}$ $=U_{t}{\rm d}t$
with unknown parameter $\theta\in\mathbb{R}$. The data $X_{t}^{\dagger}$ is generated with $\theta_{\ast}=1$, while the SGD iteration starts from $\theta_{0}=2$. The sample interval is set to $\tau=0.025$. The data is generated with a much smaller step-size of $h=\tau/100$. A total of $N=10^{5}$ positions $X_{t_{n}}$, $n=1,\ldots,N$ have been generated. The learning rate is given by $\alpha_{t_{n}}=6.0/n$. See Figure 1 middle panel for the time evolved estimates $\theta_{n}$ generated by the modified SGD (34). These estimates clearly converge to $\theta_{\ast}$. In Figure 1 we also display results from the standard SGD as well as the modified Kalman filter presented in section 4. For the Kalman filter, a Gaussian prior was used with $m_{\rm{prior}}=2$ and $\Sigma_{\rm{prior}}=6$. The diffusion constant $\sigma$ was assumed to be unknown and treated as part of the prior. While the standard SGD converges slightly faster than the modified SGD, it converges to a wrong parameter value. The estimates of the modified Kalman filter are very similar to estimates of the modified SGD. An unmodified Kalman filter (not shown) on the other hand leads to a similar estimation bias as the standard SGD.
Figure 1. Left panel: Estimated parameter value from standard SGD as a function of the iteration index. The iteration starts at $\theta_{0}=2$ and the reference value is $\theta_{\ast}=1$. A systematic bias can clearly be seen. Middle panel: Same data is used in the modified SGD. The modified SGD algorithm converges to the correct reference value. Right panel: Same data is used in the unbiased Kalman filter with prior distribution $N(2,1)$ and $\sigma$ treated as part of the prior. The lightly colored lines indicate the $1\sigma$-intervals of the Bayesian estimates.
6. Conclusions
In this note, we have proposed a simple-to-implement modification of standard SGD and Kalman filter formulations for online parameter estimation from position data only. The essential idea is to establish an Itô-like modification that re-establishes the martingale property in the presence of temporal correlations. The approach has been inspired by related work [1, 17] on multi-scale diffusion and parameter estimation for the arising reduced equations. We finally note that a corrected MLE estimator can be based on the negative log-likelihood function
(39a) $\mathcal{L}(\theta,\sigma)$ $=\frac{1}{2\sigma}\sum_{n=1}^{N-3}\theta^{\rm T}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})^{\rm T}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})\theta\tau$
(39b) $\qquad-\,\frac{1}{\sigma}\sum_{n=1}^{N-3}\theta^{\rm T}F(X_{t_{n}}^{\dagger},\tilde{U}_{n})^{\rm T}\left(\tilde{U}_{n+5/2}-\tilde{U}_{n+3/2}-f(X_{t_{n}}^{\dagger},\tilde{U}_{n})\tau\right)$
(39c) $\qquad+\,\frac{3}{4\tau\sigma}\sum_{n=1}^{N-3}\|\tilde{U}_{n+5/2}-\tilde{U}_{n+3/2}\|^{2}+\frac{N-3}{2}\log\sigma.$
Compare this likelihood function to the contrast function presented in [12], which is based on adding an appropriate correction term. Please also note that (39a)-(39b) replaces the ergodic least squares functional (3) using measured positions and estimated velocities and that (39c) delivers the estimator (33).
Future directions include the investigation of the proposed estimators in the context multi-scale diffusion when (1) arises from homogenization of an underlying multi-scale diffusion process and the inclusion of observation errors in the positions $X_{t}$. In the former case, a combination of the approaches taken here and in [17] emerges as a feasible starting point.
Acknowledgments.
This work has been partially funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294.